OpenStudy (anonymous):

For a given statement Pn, write the statements P1, Pk, and Pk+1.

OpenStudy (anonymous):

\[2+4+6+...+2n=n(n+1)\]

OpenStudy (anonymous):

I have already figured out that n=1 is true and I do not know how to solve \[2+4+6+...+2k=k(k+1)\]

OpenStudy (anonymous):

you get to assume that \[2+4+6+...+2k=k(k1) \] and use that to prove that \[2+4+6+...+2k+2(k+1)=(k+1)(k+2)\]

OpenStudy (anonymous):

sorry i mean you get to assume that \[2+4+6+...+2k=k(k+1)\]

OpenStudy (anonymous):

Oh okay so you just change form to n=k, but you do not solve it?

OpenStudy (anonymous):

replace the \(2+4+6+...+2k\) in the expression \[2+4+6+...+2k+2(k+1)\] by \(k(k+1)\) "by induction" then see if you can prove that \[k(k+1)+2(k+1)=(k+1)(k+2)\] by algebra

OpenStudy (anonymous):

wait where did the k(k+1) come from at the other side? I thought it would be \[2(k+1)=k+1(k+1+1)\]

OpenStudy (anonymous):

I'm just sitting here learning

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

have you ever done a proof by induction or is this the first?

OpenStudy (anonymous):

you get to assume the statement \(P_k\) which is what you get when you replace \(n\) by \(k\)

OpenStudy (sirm3d):

clarifications on the inductive step "assume n=k" \[2+4+6+\cdots+2k=k(k+1)\] the left hand side consists of k terms the right hand side is an expression equivalent to the sum of k terms prove true for n=k+1 the left hand side has k+1 terms, \((2+4+6+\cdots+2k)+2(k+1)\) the first k terms is enclosed in parentheses, the (k+1)st (added) term is \(2(k+1)\) the right hand side is an expression obtained by replacing n by k+1 \[(k+1)([k+1]+1)\] or \[(k+1)(k+2)\] after simplification

OpenStudy (anonymous):

in other words you get to assume that \[2+4+6+...+2k=k(k+1)\]

OpenStudy (anonymous):

whoa wait what? This is the first problem I have seen like this one, but I have done mathematical induction before

OpenStudy (anonymous):

we take that as a given

OpenStudy (anonymous):

in other words, we assume that we know that \[2+4+6+...+2k=k(k+1)\] is that part okay?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now using the above, you are trying to show that the statement \(P_{k+1}\) is true the statement \(P_{k+1}\) is what you get when you replace \(n\) by \(k+1\)

OpenStudy (anonymous):

that is, you are going to try to show that \[2+4+6+...+2k+2(k+1)=(k+1)(k+2)\] is that part okay?

OpenStudy (anonymous):

wait how did you get the right side?

OpenStudy (anonymous):

i replaced every \(n\) in the original statement by \(k+1\)

OpenStudy (anonymous):

oh okay so its k+1+1 which is k+2 I ssee that now

OpenStudy (anonymous):

it was \(n(n+1)\) if you put \(n=k+1\) you get \((k+1)(k+2)\)

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we use the "induction hypotheses" which means we get to replace \(2+4+6+...+2k\) by \(k(k+1)\)

OpenStudy (anonymous):

that is called the "inductive step" do you understand that? we use what we get to assume

OpenStudy (anonymous):

\[\overbrace{2+4+6+...+2k}+2(k+1)\] that part on the left gets replaces by \(k(k+1)\)

OpenStudy (anonymous):

okay so then its \[k(k+1)+k(k+1)=(k+1)(k+2)\]

OpenStudy (anonymous):

that is what you are going to have to show by algebra, that the left hand side is equal to the right hand side

OpenStudy (anonymous):

oh wait, there is a mistake in what you wrote

OpenStudy (anonymous):

it should be \[k(k+1)+2(k+1)=(k+1)(k+2)\]

OpenStudy (anonymous):

oops I just saw that

OpenStudy (anonymous):

\[\overbrace{2+4+6+...+2k}+2(k+1)\] \[k(k+1)+2(k+1)\]

OpenStudy (anonymous):

you want to show that this is \((k+1)(k+2)\) that is the algebra part. that is just algebra the main part is trying to figure out how to get the proof started,

OpenStudy (anonymous):

the algebra part is usually easy in this case you have \[k(k+1)+2(k+1)\] both terms have a common factor of \(k+1)\) so you can factor it and get \[(k+2)(k+1)\] in one steep

OpenStudy (anonymous):

oh okay! thank you for the help!

OpenStudy (anonymous):

yw